| ID | 
Date | 
Author | 
Category | 
Subject | 
Text | 
  | 
| 
  1  
 | 
Wed Jun 15 13:11:38 2022 | 
Guy Leckenby | DAQ | Signal Diagram | Attached is the signal mapping that will be 
used for the BEAVERTAIL test.
 
 | 
    | 
| 
  2  
 | 
Fri Jun 17 15:32:47 2022 | 
Guy Leckenby | Electronics | Preamp Settings | Bias Select: In general, higher bias resistors 
-> lower electronic noise. However, higher 
resistance = larger voltage drop due to leakage | 
    | 
| 
  3  
 | 
Mon Jun 20 10:36:56 2022 | 
Guy Leckenby | DAQ | DAQ Startup Commands | Three terminals/processes are required to 
record data on a local machine.
 
1. log in x86l-131. 
 | 
   | 
| 
  4  
 | 
Tue Jun 21 02:37:07 2022 | 
Guy Leckenby, Iris Dillmann, Chris Griffin | General | Run notes | Beam: 208Pb(82+), 269.0 MeV/u, ~2*10^7 ions 
in ESR
 
 | 
      | 
| 
  5  
 | 
Wed Jun 22 18:01:18 2022 | 
Guy Leckenby | Detectors | Detector Arrangment | Attached is the detector arrangement for the 
3 setups used during the tests. | 
    | 
| 
  6  
 | 
Wed Jun 22 19:14:12 2022 | 
Guy Leckenby | DAQ | FBEX Script Choices | Here were the f_user.c choices we implemented 
for this experiment.
 
 | 
    | 
| 
  7  
 | 
Tue Jun 28 17:42:53 2022 | 
Guy Leckenby | DAQ | Cabling Changes | In the previous signal map, I incorrectly 
mapped the DSSD channels. To correctly read 
out the DSSD for the new test, we will swap | 
    | 
| 
  8  
 | 
Fri Jul 29 15:33:08 2022 | 
Guy Leckenby | Electronics | Results of Noise Testing | Uwe and I were able to get some noise testing 
done before the ESR was closed with asbestos 
exposure. We used the new flange so didn't | 
     | 
| 
  9  
 | 
Mon Nov  7 14:14:38 2022 | 
Chris Griffin, Guy Leckenby | Electronics | MSI-8 module details | MSI-8 module details.
 
 
MMPR1 card configurations:
 | 
   | 
| 
  10  
 | 
Fri Nov 11 10:27:59 2022 | 
Chris Griffin, Guy Leckenby | Detectors | Current leakage currents | Attached are photos of leakage currents for 
the current GAGG and BGO setups.
 
This setup uses the new flange (blue gas | 
     | 
| 
  11  
 | 
Tue Nov 15 14:30:25 2022 | 
Chris Griffin | Electronics | Old flange noise | Last week we tested both the new and old flanges 
with the GAGG housing set up described in 
the previous ELog entry.
 | 
    | 
| 
  12  
 | 
Thu Nov 17 13:42:16 2022 | 
Chris Griffin | Detectors | Detector config changes | I have changed the configuration of Housing 
2 (BGO) to account for the detector being 
fixed (bond wires) and provide a "plug-and-play" | 
   | 
| 
  14  
 | 
Mon May 29 12:23:48 2023 | 
Guy Leckenby, Chris Griffin | Electronics | Old flange noise re-tests | Whilst constructing a new flange to replace 
the old one, we did some tests again to compare 
the performance of new vs old flange.
 | 
        | 
| 
  15  
 | 
Thu Jun  1 09:39:24 2023 | 
Guy Leckenby, Chris Griffin | Electronics | New new flange test | We tested the new new flange using the same 
setup as in ELOG entry 14. The noise level 
seemed to be somewhere between the middle | 
      | 
| 
  16  
 | 
Thu Jun  1 10:42:31 2023 | 
Guy Leckenby | Electronics | Summary of Preamp Testing | MSI-8 and MPR-16 preamps have nominal differential 
output ranges of +/-1.8V and +/-1.2V respectively 
at maximum input (ie 5GeV).
 | 
    | 
| 
  17  
 | 
Tue Dec 12 12:57:16 2023 | 
Jan | DAQ | test lmd-files with pulser | We took a few lmd-files with different settings 
of the febex DAQ to provide data for go4 
development, when the DAQ is not running.
 | 
   | 
| 
  18  
 | 
Mon Mar 24 14:44:14 2025 | 
C Griffin, M Cockshutt, I Dillmann | General | April 2025 experiment setup | Further PLEIADES tests are to be carried out 
during E018 and E203. This will be an important 
test of PLEIADES operation in isochronous | 
   | 
| 
  19  
 | 
Fri Mar 28 13:31:53 2025 | 
CG, ID, MC, JG | General | Dual PLEIADES installation at ESR | PLEIADES detectors were installed in both 
the inside and outside pockets after the 
dipole behind the gas jet.
 | 
       | 
| 
  20  
 | 
Fri Mar 28 20:45:08 2025 | 
 CG | Electronics | MSI-8 configurations | The MSI-8 modules were reconfigured to properly 
instrument the current setup.
 
 | 
        | 
| 
  21  
 | 
Fri Mar 28 22:45:24 2025 | 
CG | Detectors | Bias problems | After testing the biasing scheme for all detectors, 
two problems emerge.
 
 | 
     |